OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 5x + 3y = 8 y = −2x + 1

OpenStudy (anonymous):

i know how to do elimination and substitution but not fully

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

no

OpenStudy (anonymous):

help @tcarroll010

OpenStudy (anonymous):

Choosing substitution as your choice for solving is the right way to go here because that second equation is already in a form that specifies the variable because you have: y equals "something on the right" You can take that "something on the right" from equation #2 and put that in the first equation wherever you have "y". That way you can solve equation #1 as it will now be in only one variable, "x".

OpenStudy (anonymous):

ok so it would be 3y=8-5x for the first part which equals y?

OpenStudy (anonymous):

and then u plug that into the equation?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

equation #2

OpenStudy (anonymous):

yes. so then the equation would be 8-5x −2x + 1??

OpenStudy (anonymous):

or 3(8-5x) −2x + 1??

OpenStudy (anonymous):

You're not following the steps I outlined. What does y equal from equation #2?

OpenStudy (anonymous):

y=-2x+1

OpenStudy (anonymous):

im not really good at this sorry

OpenStudy (anonymous):

The equations have a left side and a right side. I said to take that right side of equation #2 and put that (just the right side, because that is another form of "y") into equation #1 where you see the "y".

OpenStudy (anonymous):

oh ok so its 5x + 3(-2x+1) = 8

OpenStudy (anonymous):

Yes, now you're getting it! That's the "substitution" concept. Now you just simplify that and solve for "x". When you get "x", you can go back to either ORIGINAL equation, put that "x" in, and solve for "y".

OpenStudy (anonymous):

oh ok so when i simplify it it would be 5x + 3 (-2x+1)=8 5x -6x + 3 = 8 -1x + 3= 8 - 3 - 3 -1x = 5 x = -5

OpenStudy (anonymous):

Yes, that's "x", so just solve for "y" now, using what I pointed out.

OpenStudy (anonymous):

and then 5 (-5) + 3y = 8 -10 + 3y = 8 +10 +10 3y = 18 3y/3 = 18/3 y = 6

OpenStudy (anonymous):

5(-5) is not -10

OpenStudy (anonymous):

so y = 11

OpenStudy (anonymous):

sorry i wasnt thinkinh

OpenStudy (anonymous):

Yes, you got it right, good job!

OpenStudy (anonymous):

All good now, @awk0perf ?

OpenStudy (anonymous):

yes ty

OpenStudy (anonymous):

uw! Nice working with you!

OpenStudy (anonymous):

thanks